You want to create a 0.25 M Sodium Bicarbonate solution. You weigh 5.00 grams of Sodium Bicarbonate. How much water is needed?


1
Expert's answer
2022-02-22T08:49:01-0500

Molarity (M) = no. of moles "\/" volume

molar mass of sodium bicarbonate "NaHCO_3=84"

No. of moles of sodium bicarbonate "=\frac{5}{84}=0.05952"

volume of water needed = "\frac{no. \space of \space moles}{molarity}= \frac{0.25}{0.05952}=4.2litres"